OpenStudy (anonymous):

Will give Medal! 1. The membership in the math club at Rogers Junior High increased by 4% from last year to this year. The president of the math club used the following expression to find the number of members in the club this year. m + m(0.04) Which of the following is another expression that could be used to get the same results? m + 4 104m 1.04m m + 1.04

OpenStudy (amanda0618):

Wait...Is there an error in your question? I feel like something is wrong.

OpenStudy (anonymous):

No I don't think so

OpenStudy (amanda0618):

hmm

OpenStudy (anonymous):

I copied it of of the website

OpenStudy (cwrw238):

nothing wrong with it

OpenStudy (anonymous):

ok cool

OpenStudy (cwrw238):

4% = 0.04

OpenStudy (cwrw238):

104% = 1.04

OpenStudy (cwrw238):

that should help

OpenStudy (cwrw238):

let m = 100%

OpenStudy (anonymous):

Still don't get it

OpenStudy (texaschic101):

m + m (0.04) ..... keep in mind that m is the same as 1m. Now distribute through the parenthesis. 1m + .04m --- combine like terms and what do you get ?

OpenStudy (anonymous):

1.04m

OpenStudy (texaschic101):

yep

OpenStudy (anonymous):

Awesome Sauce!

OpenStudy (texaschic101):

lol :)

OpenStudy (anonymous):

I have another one if you don't mind...

OpenStudy (texaschic101):

I can try...can't make no promises

OpenStudy (anonymous):

Ok cool...Simplify the following expression. (41x + 62y) + (-20x + 49y) 61x + 111y 21x + 13y 61x + 13y 21x + 111y

OpenStudy (texaschic101):

41x + 62y - 20x + 49y -- (I distributed the + through the parenthesis)-- combine like terms 21x + 111y

OpenStudy (anonymous):

Awesome! I have another one!

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

Simplify the following expression. (48x + 59y) + (45x - 13y) 107x + 72y 93x + 46y 107x + 46y 93x + 72y

OpenStudy (texaschic101):

same thing as the last problem...distribute the + through the parenthesis and you get : 48x + 59y + 45x - 13y -- combine like terms your turn... so what is 48x + 45x ? and what is 59y - 13y ?

OpenStudy (texaschic101):

any thoughts ?

OpenStudy (anonymous):

93x and 46y

OpenStudy (texaschic101):

awesome sauce....good job...so the answer is : 93x + 46y

OpenStudy (anonymous):

LOL!!! Awesome Sauce!

OpenStudy (texaschic101):

you deserve a medal :)

OpenStudy (anonymous):

Awwww Thank you!

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

I have another and another and more so I hope you don't get tired of me... LOL!

OpenStudy (texaschic101):

I am ready....whatcha got ?

OpenStudy (anonymous):

Okey dokey Francis bought a new dress that was discounted by 29%. She used the following expression to find the price of the dress after the discount was applied. d - (0.29)d Which of the following is another expression that could be used to get the same results? d - 29 71d d - 0.29 0.71d

OpenStudy (texaschic101):

1d - (0.29)d = 1d - 0.29d = 0.71d

OpenStudy (anonymous):

Thank you! Simplify the following expression. (18x - 11y) + (41x + 62y) 7x + 51y 59x + 73y 7x + 73y 59x + 51y

OpenStudy (texaschic101):

just remember that when you just have a variable (a letter)...if there is no number in front, it is understood to be 1

OpenStudy (anonymous):

ok I will remember.

OpenStudy (texaschic101):

18x - 11y + 41x + 62y -- combine like terms 59x + 51y

OpenStudy (anonymous):

Awesome Sauce! Simplify the following expression. (15x + 6y) + (50x + 35y) 21x + 50y 65x + 50y 21x + 41y 65x + 41y

OpenStudy (texaschic101):

15x + 6y + 50x + 35y -- combine like terms 65x + 41y

OpenStudy (anonymous):

Thank you! Which algebraic expression is equivalent to the expression below? 12(3x + 4) - 7x 29x + 48 36x + 41 43x + 48 29x - 4

OpenStudy (texaschic101):

12(3x + 4) - 7x --- distribute through the parenthesis 12 * 3x + 12 * 4 - 7x 36x + 48 - 7x -- combine like terms 29x + 48

OpenStudy (anonymous):

Thanks!Oh and I like your profile pic! Which of the following is equivalent to the expression below? (3x + 9) + (17 - 3x) 6x + 26 26 8 3x + 26

OpenStudy (anonymous):

@texaschic101 where did you go?

OpenStudy (texaschic101):

3x + 9 + 17 - 3x -- combine like terms 26 ...because the 3x's cancel out and we are left with 9 + 17 which equals 26

OpenStudy (texaschic101):

I sometimes have to refresh my page...my computer is not that good. I am not leaving

OpenStudy (anonymous):

Awesome! I have to do that sometimes too. Which algebraic expression is equivalent to the expression below? 9(2x - 4) + 9x 27x - 4 9x + 36 27x - 36 9x - 36

OpenStudy (texaschic101):

9(2x - 4) + 9x -- distribute through the parenthesis 9 * 2x + 9 * -4 + 9x 18x - 36 + 9x -- combine like terms 27x - 36

OpenStudy (anonymous):

Ok this will be a bit diffrent then what we have been doing. So be prepared to be amazed.

OpenStudy (texaschic101):

lol...amaze me :)

OpenStudy (anonymous):

Ok don't say I didn't warn you! LOL! Directions: Type the correct answer in the box. Use numerals instead of words. Janette went to Tire Mart and paid $122.57 per tire for 4 tires and a flat rate fee for installation for a total of $555.28 before taxes. Monique went to Fast Wheels and paid $136.71 per tire for 4 tires and a flat rate fee for installation for a total of $585.34 before taxes. The difference in the amount that they paid in installation charges is $.

OpenStudy (texaschic101):

Tire Mart 122.57 per tire for 4 tires....so thats 122.57 * 4 = 490.28....and if she ended up paying 555.28, then the installation costs 555.28 - 490.28 = 65 installation. Fast wheels 136.71 per tire for 4 tires...so thats 136.71 * 4 = 546.84..and if she ended up paying 585.34, then the installation costs : 585.34 - 546.84 = 38.50 installation. SO the difference is : 65 - 38.50 = 26.50

OpenStudy (anonymous):

So it's 26.50?

OpenStudy (texaschic101):

yes...thats the difference Tire mart charges 65 installation and fast wheels charges 38.50 installation...for a difference of 26.50

OpenStudy (anonymous):

YaY! You are such a big help! The perimeter of a rectangle is 46 inches. If the length of the rectangle is 14 inches, which equation could be used to find the width, x? x + 2(14) = 46 x + 14 = 46 14(x + 2) = 46 2(x + 14) = 46

OpenStudy (texaschic101):

formula for perimeter of rectangle : P = 2(L + W) P = 46 L = 14 W = x so we sub 46 = 2(14 + x)...rearrange it .... 2(x + 14) = 46

OpenStudy (anonymous):

Your Awesome! Susan's math teacher assigned x homework problems on Monday. On Tuesday, she assigned 14 more problems. Over the two days, a total of 22 homework problems were assigned. Which equation could be used to find x, the number of problems assigned on Monday? x + 14 = 22 22 + x = 14 x - 14 = 22 14x = 22

OpenStudy (texaschic101):

x problems on Monday....she was then assigned 14 more problems....x + 14...for a total of 22....so your equation is : x + 14 = 22

OpenStudy (anonymous):

Cool! Directions: Select all the correct answers. Which of the situations correspond to the equation below? Chase donated t to the school's fundraiser, which put their fundraising total at $339.50 or 35% of their goal target of $850. After paying t toward her tuition, Julie still owed 35% of her total tuition of $850.00, leaving an amount due of $339.50. Manuel received a 35% discount for a savings of $339.50 by purchasing a set of four tires at $212.50 each plus a road hazard warranty for t. Monica read in the paper that the Formal Gown Warehouse is offering 35% of the cost of a gown toward a pair of dress shoes. The gown that Monica wants cost $850.00 and the shoes that go with the dress are also on sale for t, which would give her a total savings of $339.50. Rainy's neighbor was having a new fence installed and asked her to pay 35% of the cost based on the portion of the fence that ran between their houses. He asked her for $339.50 to cover $850.00 worth of materials plus labor costs, t.

OpenStudy (texaschic101):

well crap...I am not sure. I want to say the third one...you might need someone else to look at this one

OpenStudy (anonymous):

ok I will look for someone who helped me before. Stay here I will be back

OpenStudy (texaschic101):

wait...it might be 4th one

OpenStudy (anonymous):

Are you sure?

OpenStudy (texaschic101):

not 100%...sorry...I am leaning towards the 4th one. You want to try to find the person who helped you before ?...just to make sure

OpenStudy (anonymous):

well It is which ones that will fit it. So multiple choice. But I will get them.

OpenStudy (texaschic101):

I am sorry but I have to go now. I will be back on later. I hope I was able to help you...even just a little bit. I wish you luck with all this stuff :)

OpenStudy (anonymous):

Oh ok well I am about to go to got a appointment. But thank you so much for all of your help! Thank you!

OpenStudy (texaschic101):

YW :) AWESOME SAUCE....lol

OpenStudy (anonymous):

LOL! Awesome sauce to you too! Bye!

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@.Sam. please help!

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@monsterfox1997

OpenStudy (anonymous):

@TheRaggedyDoctor

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Amy and Stephanie were playing a number game. Amy picked a number and told Stephanie that if she added 8 to the number and multiplied the sum by 3, it would equal 9. Which equation could Stephanie use to find the number, x? 3 + (x + 8) = 9 8(x + 3) = 9 3(x) + 8 = 9 3(x + 8) = 9

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

@TheRaggedyDoctor Sorry but I need help

OpenStudy (theraggedydoctor):

OOPS SORRY

OpenStudy (theraggedydoctor):

Okay so.. add 8 to the number and multiply the sum of the two numbers by 3. Which do you think it is?

OpenStudy (anonymous):

It is ok but never mind I have to g. I will be here later are you?

OpenStudy (theraggedydoctor):

I might be, lol

OpenStudy (anonymous):

Ok thank you very much! Bye!